OpenStudy (maheshmeghwal9):

If \[z, \omega \ne 0.\] & \[|z|= |\omega|.\] and \[\arg.(z)+\arg.(\omega)=\pi \space \space then \space \space z=...?\]

OpenStudy (maheshmeghwal9):

@satellite73 Please help:)

OpenStudy (anonymous):

seems like we have an infinite number of possibilities am i missing something?

OpenStudy (anonymous):

\[z=re^{i\theta}\] \[\omega=re^{i(\pi-\theta)}\]

OpenStudy (anonymous):

\[\omega=r(\cos(\theta)+i\sin(\theta))\] \[z=r(\cos(\pi-\theta)+i\sin(\pi-\theta))\] \[z=r(\sin(\theta)+i\cos(\theta))\] by the cofunction identity, so if \[\omega=a+b\] then \[z=b+ai\]

OpenStudy (anonymous):

@maheshmeghwal9 \[z=\left| z \right| e^{i \arg(z)}=\left| \omega \right| e^{i (\pi-\arg(\omega))}=\left| \omega \right|e^{i \pi}e^{-i \arg(\omega)}=-\left| \omega \right|e^{-i \arg(\omega)}=-\bar{\omega}\] since \[e^{i \pi}=-1\]

OpenStudy (maheshmeghwal9):

@mukushla I have a doubt. How did u write \[\huge{-|\omega|e^{-i \arg(\omega)}=- \bar \omega.}\]

OpenStudy (maheshmeghwal9):

i mean why.

OpenStudy (anonymous):

\[\left| \omega \right|e^{-i \arg ( \omega)}=\left| \omega \right| (\cos (\arg ( \omega))-i \sin (\arg ( \omega))=\bar{\omega} \\ since\\ \omega=\left| \omega \right| (\cos (\arg ( \omega))+i \sin (\arg ( \omega))\]

OpenStudy (maheshmeghwal9):

oh i see thanx a lot i gt it:)